OpenStudy (anonymous):

how do I convert 575m^2 into ha?

OpenStudy (anonymous):

i can find the area but i don't understand the ha part?

OpenStudy (anonymous):

is one ha 10,000?

OpenStudy (anonymous):

1 acre = .40 hectacre (HA)

OpenStudy (anonymous):

how many yards are in an acre?

OpenStudy (anonymous):

it has to be in decimal form. I thought I just divided.

OpenStudy (anonymous):

no you have to convert that to HA first using m2 to yd to acre then to HA

OpenStudy (anonymous):

1 sq yard = .84 sq meters

OpenStudy (anonymous):

the question is .. a rectangular field is 25m by 23 m. find the area in m^2 (which I got 575m^2) and in ha.

OpenStudy (anonymous):

1 ha = 10000 (m^2) x- 575 (m^2) 10000x = 575 x= 575/10000 x= 0.0575

OpenStudy (anonymous):

69.57 yd = 1acre

OpenStudy (anonymous):

of course x= 0.0575 ha

OpenStudy (anonymous):

that's what i was thinking natka but it told me it was wrong.thats why im lost

OpenStudy (anonymous):

maybe there is a mistake in your answer book? Sometimes it happens...

OpenStudy (anonymous):

I got 3.93 HA

OpenStudy (anonymous):

its 0.06 lol

OpenStudy (anonymous):

i get to do a new problem

OpenStudy (anonymous):

oh :)

OpenStudy (anonymous):

i hve not clue how they got that. the example doesn't show me much

OpenStudy (anonymous):

now i have to convert 598m^2 lol